OpenStudy (anonymous):

order of operations: 2/3 + 1/5 -------- 1/2 i got 26/15, but im not sure its quite right

OpenStudy (anonymous):

13/30

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

26/15 your right

OpenStudy (anonymous):

2/3 + 1/5 = 10/15 + 3/15 = 13/15 ----- (13/15) / (1/2) = 13/15 x 2 =(13 x 2)/15

OpenStudy (anonymous):

so think of the numerator as having parentheses around it, and then consider the denominator, so following the math i posted above, u get 26/15, you were right!

OpenStudy (anonymous):

26/15 ^^

OpenStudy (anonymous):

thank youuuu!